OpenStudy (anonymous):

6(3a+1)-30=3(2a-4) ...?

OpenStudy (anonymous):

Got 4 but when I plugged it back in it did not equal.

OpenStudy (anonymous):

a = 1...

OpenStudy (anonymous):

Can you show work...thanks

OpenStudy (anonymous):

18a+6-30=6a-12 12a-24=-12 12a =12 a=1

OpenStudy (anonymous):

Oh...I was trying to divide 6a from 18a...you combine by addition or subtraction the 6a and 12a?

OpenStudy (anonymous):

I subtracted 6a from both sides.

OpenStudy (anonymous):

I mean 18 a...sorry

OpenStudy (anonymous):

Gotcha thanks....

OpenStudy (anonymous):

6(3a+1)-30=3(2a-4)